OpenStudy (anonymous):

Simplify 2(3x + 4y) - 2(x - 2y) A: 4x + 2y B: 4x + 8y C: 4x + 12y D: 8x + 8y

OpenStudy (anonymous):

You can expand the brackets by multiplying each thing inside the brackets by the thing on the outside. For example, the first bracket is: \[2(3x+4y)=2 \times 3x + 2 \times 4y = 6x + y\]

OpenStudy (anonymous):

Sorry that last line should finish 6x + 8y.